OpenStudy (lifeisadangerousgame):

Could someone explain why this is (x+2)?

OpenStudy (lifeisadangerousgame):

I don't understand why the answer is (x+2)e^x

OpenStudy (3mar):

May I help?

OpenStudy (lifeisadangerousgame):

Sure!

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Are you familiar with the derivatives?

OpenStudy (lifeisadangerousgame):

Yes

OpenStudy (lifeisadangerousgame):

I'm just starting to learn the product rule

OpenStudy (3mar):

at f'(x) or at f''(x)??

OpenStudy (lifeisadangerousgame):

at f(x)

OpenStudy (3mar):

mean f'(x)

OpenStudy (lifeisadangerousgame):

What?

OpenStudy (3mar):

Where are you stuck?

OpenStudy (lifeisadangerousgame):

I'm stuck on why the end answer is (x+2)e^x. Why x + 2?

OpenStudy (3mar):

ok the result of f'(x) =(x+1)e^x got it? satisfied?

OpenStudy (lifeisadangerousgame):

I know that, I'm talking about part b)

OpenStudy (3mar):

Ok I just want to be agree about one point where we will take off!

OpenStudy (3mar):

f''(x)=(x+1).[e^x]+(e^x)[1]

OpenStudy (3mar):

ok here?

OpenStudy (lifeisadangerousgame):

Yeah

OpenStudy (3mar):

So can you simplify it a little bit?

OpenStudy (lifeisadangerousgame):

(x+1)e^x + e^x

OpenStudy (3mar):

take e^x as a common!

OpenStudy (3mar):

@LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

I don't really understand, but that's okay! Thank yo ufor taking the time to help me

OpenStudy (3mar):

Please can I continue?

OpenStudy (lifeisadangerousgame):

If you want you can, but I kind of already started working on the quotient rule

OpenStudy (3mar):

One min , no more! May I?

OpenStudy (3mar):

\[(x+1)e^x + e^x=e^x[(x+1)+1]=e^x[x+2]\] That is it! I hope you got what you are looking for!

OpenStudy (lifeisadangerousgame):

Ohh! That makes sense. Thank you!!

OpenStudy (3mar):

I told you it takes 1 min no more Hope that helps